# Deep Learning Improves Photometric Redshifts in All Regions of Color Space

### v1

@@ 1 Introduction
Next-generation imaging surveys such as the Euclid mission (Laureijs et al., 2011), the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST; Ivezić et al., 2019), and the Nancy Grace Roman Space Telescope (Roman; Spergel et al., 2015; Akeson et al., 2019) will obtain deep multi-band optical and near-infrared imaging over thousands of square degrees. Nearly all of the extragalactic science drivers of these surveys, such as cosmology, galaxy evolution, and transients, will require redshift estimates (Newman & Gruen, 2022). Ideally, spectroscopic redshifts (spec-’s) would be used to estimate distances, but obtaining spectra for any significant fraction of the photometrically-detected objects is completely infeasible with current facilities (e.g., Newman et al., 2015). Instead, these imaging surveys must rely on photometric redshifts (photo-’s; e.g., The LSST Dark Energy Science Collaboration et al. 2018).
@@
Photometry-based supervised ML algorithms have produced more precise photo- estimates for imaging surveys than physically-motivated techniques, such as template fitting (which involves fitting galaxy spectral templates to a galaxy’s observed integrated photometry), when sufficient training data is available (Salvato et al., 2019). More recently, the advent of convolutional neural networks (CNNs) has enabled deep learning photo- models that can leverage pixel-level color information from the images directly (see Huertas-Company & Lanusse 2023 for a review of deep learning methods applied to galaxy surveys). These modern imaged-based deep neural networks produce the current best photo- predictions if given training samples with complete coverage of color–magnitude–redshift space, such as the Sloan Digital Sky Survey (SDSS), even with training sets that are relatively small (e.g., Pasquet et al., 2019; Hayat et al., 2021; Henghes et al., 2022; Dey et al., 2022a; Treyer et al., 2024; Ait Ouahmed et al., 2024). Despite these success, it remains unclear exactly why they perform better than classical ML methods on global metrics, which motivates this work.
@@
Strong performance on global metrics can mask offsetting systematic errors in local regions of input parameter space. An extreme case of this effect is the trainZ photo- estimator from the LSST Photo- Data Challenge (Schmidt et al., 2020), which simply predicted the redshift distribution of the training set as the photo- probability distribution function for all objects, regardless of their photometry. By construction, the trainZ estimator achieved the ideal distribution of the probability integral transform when considered globally, beating out the other 11 widely-used photo- methods. However, it performed the worst in local regions of color space (as characterized by the conditional density estimation loss; see, e.g., Izbicki & Lee 2017). A full conditional assessment by regressing across color space (e.g., using the Cal-PIT method of Dey et al. 2022b) is beyond the scope of this paper. Rather, we simply partition color space using a self-organizing map (SOM; Kohonen 1982, 2001; see Section 4) and assess photo- performance in neighborhoods of color space.
@@ 2 Data
Our analysis uses the SDSS (Gunn et al., 1998, 2006; York et al., 2000; Smee et al., 2013) Main Galaxy Sample (MGS; Strauss et al. 2002) from Data Release 12 (DR12; Alam et al. 2015). We started with an identical selection criteria as Pasquet et al. (2019); namely, we selected galaxies based on their dereddened -band Petrosian magnitude () and spec- (), which yielded 516,525 galaxies. We split this sample into the same training (464,873 galaxies) and test (51,652 galaxies) sets as Dey et al. (2022a). We further removed galaxies with very faint dereddened model magnitudes (, , , , or ) to ensure reliable photometry. These cuts reduced our dataset to 509,731 galaxies, with 458,740 galaxies in our training set and 50,991 galaxies in our test set.
We also used galaxy properties from the SDSS photometric catalog (Stoughton et al., 2002), such as fracDev_r (the weight of the de Vaucouleurs component in a combined de Vaucouleurs + exponential profile fit in the -band) and expAB_r (the semi-minor/semi-major axis ratio for the exponential profile fit in -band, which we note is only reliable for spirals), which we adopt as the bulge-to-total ratio and inclination, respectively. Throughout this work, we use the bulge-to-total ratio as a single parameter to summarize a galaxy’s rest-frame color (leveraging the well-known trend between morphology and rest-frame color for low- galaxies; e.g., Blanton et al. 2003), instead of the specific star formation rate, which requires highly accurate spectra that not all surveys can obtain, including SDSS.
@@ 3 Photometric Redshifts
To compare the differences between photometry-based classical ML photo- methods and image-based deep learning photo- methods, we considered two approaches in each category. For classical ML methods, we adopted the photo-’s from the SDSS DR12 catalog as estimated by Beck et al. (2016, hereafter B16) and our own photo-’s determined using a random forest. For the deep learning models, we investigated photo-’s from Dey et al. (2022a) and Pasquet et al. (2019).
@@ 3.1 SDSS DR12 Catalog
For the SDSS DR12 catalog photo-’s, B16 used local linear regression, which fits a first-order polynomial to a small number of neighbors with labels (i.e., spec-’s) in the input space (modelMag colors and cModelMag magnitude). Performing this calculation locally preserves the small-scale structure in the data and allows it to track subtle trends. B16 used 100 neighbors in their fitting procedure, though they implemented more nuanced criteria to handle under-sampled regions of color–magnitude space and to iteratively filter redshift outliers.
B16 trained their model using 1,976,978 galaxies with spec-’s from the entire SDSS DR12 spectroscopic sample—specifically, the MGS, the luminous red galaxy sample (Eisenstein et al., 2001), and the Baryon Oscillation Spectroscopic Survey (BOSS; Dawson et al., 2013)—plus additional spectroscopy from other surveys, such as DEEP2 (Newman et al., 2013), VIPERS (Garilli et al., 2014; Guzzo et al., 2014), and zCOSMOS (Lilly et al., 2009). B16 estimated photo-’s for the 208,474,076 galaxies of the SDSS primary photometric catalog, thus their aims were broader than those of the other approaches we consider, which only used MGS galaxies for training and testing. That being said, MGS galaxies dominate the information content within the color–magnitude–redshift space covered by MGS galaxies, which is further prioritized by the local nature of their regression method.
@@ 3.2 Random Forest
Inspired by previous random forest photo- efforts (e.g., Carliles et al., 2010; Carrasco Kind & Brunner, 2013; Zhou et al., 2019), we trained a random forest (Breiman, 2001) for photo- estimation using the RandomForestRegressor method from the Python package Scikit-learn (Pedregosa et al., 2011) as a classical ML alternative to the local linear regression method of B16. Random forest regression is a simple supervised learning algorithm consisting of a “forest” of decision trees that are each trained to predict a value with the final result of the algorithm being the average of each tree’s predicted value. To train the algorithm, we used the training set described in Section 2. We tested providing the random forest with multiple different sets of input features, starting with the features B16 used to train their network, as discussed in Section 3.1. To test the impact of input feature choices, we ran random forest models with several combinations of input features:
@@
Ultimately, we found only minor differences in the photo-’s by adding more features, so we chose model 1 (the same features as B16) as our fiducial random forest photo- model. However, we adopt the B16 photo-’s as our fiducial classical ML photo- method because it produced qualitatively similar results to the random forest approach, but the B16 photo-’s are widely used and are a standard comparison sample for photo- analyses.
@@ 3.3 Deep Learning Models
Despite their architectural differences, a variety of modern deep learning models—including CNNs (Pasquet et al., 2019), semi-supervised representation learning (Hayat et al., 2021), and deep capsule networks (Dey et al., 2022a)—have all significantly out-performed classical ML photo- methods (namely B16) on the SDSS MGS. In this work, we considered photo-’s from Pasquet et al. (2019) and Dey et al. (2022a), both of which were trained using the dataset described in Section 2. We adopt the latter as our fiducial deep learning photo- model, though we found very similar results with the former.
@@ 3.3.1 Capsule Networks
Hinton et al. (2011) proposed the idea of capsule networks as an alternative to CNNs, where the artificial neurons of a capsule network are organized into groups (capsules), producing the vector analog to the scalar neurons in CNNs such that they are viewpoint equivariant (as opposed to the viewpoint invariance of CNNs) and therefore could be more efficient to train. However, actually training capsule networks proved challenging until the introduction of the dynamic routing algorithm by Sabour et al. (2017) that dramatically improved the viability of their training. Soon after, Katebi et al. (2019) became the first work to apply capsule networks to galaxy images for galaxy morphology prediction.
Dey et al. (2022a) utilized a deep capsule network for their encapzulate model (compared to the shallow network proposed by Sabour et al. 2017) developed by Rajasegaran et al. (2019). Its architecture consists of three sub-networks: a deep capsule network for encoding the galaxy images to a low dimensional embedding space, a class-independent decoder network for morphological prediction that acts as a regularizer (using disk/spheroid labels from Galaxy Zoo 1; Lintott et al. 2011), and a redshift prediction network.
@@ 3.3.2 Global Photo-z Performance
Figure 1 shows versus for the B16 and encapzulate photo-’s of SDSS MGS test set galaxies. Compared to the B16 photo-’s, the encapzulate photo-’s have a smaller scatter as measured using the normalized median absolute deviation:
where . In the case of photo-’s, we define as (, ). The encapzulate photo-’s also have a significantly lower fraction of catastrophic outliers (as defined by , where ) and a smaller global bias () than the B16 photo-’s. These are generic results of comparing classical ML and deep learning photo- methods for the SDSS dataset, though it is not understood how the deep learning models leverage additional redshift information from the images.
Furthermore, there is a slight but perceptible trend of the B16 photo-’s being biased towards the mean redshift of the dataset, such that they are overestimated at low and underestimated at high , which is a common bias for ML methods called attenuation bias (see Section 6.1). Below, we compare classical ML and deep learning photo-’s for sub-samples in small neighborhoods in color space to understand how local trends impact global metrics.
@@ 4 Partitioning Color Space with a Self-Organizing Map
To study the impact of attenuation bias in local neighborhoods of color space, we partition color space using a SOM. A SOM is an unsupervised dimensionality reduction algorithm that works by grouping together objects with similar properties into discrete cells, with objects in a cell having properties that are most similar to other objects in the same cell and neighboring cells usually containing galaxies that are broadly similar. Since SOMs produce easily interpretable results with clear visualizations, they have been used for photo- estimation (e.g., Carrasco Kind & Brunner 2014; Myles et al. 2021; Wright et al. 2025) and characterizing the coverage of color space by objects with spec-’s (e.g., Masters et al. 2015, 2017; Stanford et al. 2021; Gruen et al. 2023; McCullough et al. 2024).
For our implementation, we used the minisom python package (Vettigli, 2018) to produce a two-dimensional SOM with 10 12 cells using the , , , and colors of all objects in our sample (i.e., including objects in both the training and test sets). Figure 2 shows the SOM color-coded by each cell’s mean spec- (left panel) and mean bulge-to-total ratio (B/T; right panel). Roughly speaking, mean spec- increases from left-to-right and B/T increases from bottom-to-top.
As representatives of the two main galaxy populations, we chose Cell A (8, 9), which contains red elliptical galaxies (high B/T), and Cell B (1, 3), which contains blue disks (low B/T). The other two cells of interest show more unusual scenarios. Cell C (3, 4) has an intermediate B/T and is located in the transition region between low and high B/T. It contains a diverse collection of objects, including red disk galaxies (Masters et al., 2010), post-starburst galaxies (Dressler & Gunn, 1983), and galaxies with active galactic nuclei, determined by examining the SDSS and DESI-LS images and SDSS spectra of galaxies in this cell. Finally, Cell D (0, 0) contains very low- galaxies (), a regime where the B16 algorithm suffered from a large fraction of catastrophic outliers (Geha et al., 2017) but deep learning methods performed well (e.g., Wu et al., 2022; Darragh-Ford et al., 2023).
@@ 5.1 Photo-z Scatter in Individual SOM Cells
Figure 3 shows the ratio between the values of the B16 and encapzulate photo-’s for individual SOM cells (gray points) as a function of the cell’s B/T. For all of the SOM cells, the for the B16 photo-’s is at least as large as that of encapzulate with most of the cells having an value – times larger. The binned median trend (solid blue line) shows that there is a mostly flat but slightly decreasing trend from B/T with an ratio of about –, which is similar to the global ratio of 1.5. The highest B/T cells typically have smaller from encapzulate, though a few individual cells (like Cell A) approach parity.
We found similar results when using the random forest photo-’s in lieu of the B16 photo-’s and when using the Pasquet et al. (2019) photo-’s instead of the encapzulate photo-’s. This suggests that the deep learning methods are harnessing additional redshift information from the spatially-resolved images that purely photometry-based ML methods cannot access, particularly for SOM cells with low and intermediate B/T.

### v2

@@ 1 Introduction
Next-generation imaging surveys such as the Euclid mission (euclid), the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST; ivezic2019), and the Nancy Grace Roman Space Telescope (Roman; spergel2015; akeson2019) will obtain deep multi-band optical and near-infrared imaging over thousands of square degrees. Nearly all of the extragalactic science drivers of these surveys, such as cosmology, galaxy evolution, and transients, will require redshift estimates (NewmanGruen2022). Ideally, spectroscopic redshifts (spec-’s) would be used to estimate distances, but obtaining spectra for any significant fraction of the photometrically-detected objects is completely infeasible with current facilities (e.g., Newman2015). Instead, these imaging surveys must rely on photometric redshifts (photo-’s; e.g., LsstSRD).
@@
Photometry-based supervised ML algorithms have produced more precise photo- estimates for imaging surveys than physically-motivated techniques, such as template fitting (which involves fitting galaxy spectral templates to a galaxy’s observed integrated photometry), when sufficient training data is available (Salvato2019). More recently, the advent of convolutional neural networks (CNNs) has enabled deep learning photo- models that can leverage pixel-level color information from the images directly (see HuertasCompany2023 for a review of deep learning methods applied to galaxy surveys). These modern image-based deep neural networks produce the current best photo- predictions if given training samples with complete coverage of color–magnitude–redshift space, such as the Sloan Digital Sky Survey (SDSS), even with training sets that are relatively small (e.g., pasquet2019; hayat2021; Henghes2022; dey2022; treyer2024; AitOuahmed2024). Despite these successes, it remains unclear what is the dominant mechanism driving their better performance than classical ML methods on global metrics, though previous works have claimed that deep learning algorithms are leveraging morphological information, inclination, and the presence of fainter neighboring galaxies to achieve better results (e.g., Menou2019; pasquet2019).
@@
Strong performance on global metrics can mask offsetting systematic errors in local regions of input parameter space. An extreme case of this effect is the trainZ photo- estimator from the LSST Photo- Data Challenge (Schmidt2020), which simply predicted the redshift distribution of the training set as the photo- probability distribution function for all objects, regardless of their photometry. By construction, the trainZ estimator achieved the ideal distribution of the probability integral transform when considered globally, beating out the other 11 widely-used photo- methods. However, it performed the worst in local regions of color space (as characterized by the conditional density estimation loss; see, e.g., Izbicki2017). A full conditional assessment by regressing across color space (e.g., using the Cal-PIT method of calpit) is beyond the scope of this paper. Rather, we simply partition color space using a self-organizing map (SOM; Kohonen1982; Kohonen2001; see Section 4) and assess photo- performance in neighborhoods of color space (see Jalan2024 for a similar approach, albeit with different goals, conducted contemporaneously but independently).
@@ 2 Data
Our analysis uses the SDSS (Gunn1998SDSSCamera; Gunn2006SDSSTelescope; YorkEtal2000Sdss; Smee2013SDSSSpectrograph) Main Galaxy Sample (MGS; StraussEtal2002SdssMGS) from Data Release 12 (DR12; AlamEtal2015SdssDR12). We started with an identical selection criteria as pasquet2019; namely, we selected galaxies based on their dereddened -band Petrosian magnitude () and spec- (), which yielded 516,525 galaxies. We split this sample into the same training (464,873 galaxies) and test (51,652 galaxies) sets as dey2022. We further removed galaxies with very faint dereddened model magnitudes (, , , , or ) to ensure reliable photometry. These cuts reduced our dataset to 509,731 galaxies, with 458,740 galaxies in our training set and 50,991 galaxies in our test set.
We also used galaxy properties from the SDSS photometric catalog (stoughton2002), such as fracDev_r (the weight of the de Vaucouleurs component in a combined de Vaucouleurs + exponential profile fit in the -band) and expAB_r (the semi-minor/semi-major axis ratio for the exponential profile fit in -band, which we note is only reliable for spirals), which we adopt as the bulge-to-total ratio and inclination, respectively. Throughout this work, we use the bulge-to-total ratio as a single parameter to summarize a galaxy’s rest-frame color (leveraging the well-known trend between morphology and rest-frame color for low- galaxies; e.g., blanton2003), instead of the specific star formation rate, which requires highly accurate spectra that not all surveys can obtain, including SDSS.
@@ 3 Photometric Redshifts
To compare the differences between photometry-based classical ML photo- methods and image-based deep learning photo- methods, we considered two approaches in each category. For classical ML methods, we adopted the photo-’s from the SDSS DR12 catalog as estimated by beck2016 and our own photo-’s determined using a random forest. For the deep learning models, we investigated photo-’s from dey2022 and pasquet2019.
@@ 3.1 SDSS DR12 Catalog
For the SDSS DR12 catalog photo-’s, beck2016 used local linear regression, which fits a first-order polynomial to a small number of neighbors with labels (i.e., spec-’s) in the input space (modelMag colors and cModelMag magnitude). Performing this calculation locally preserves the small-scale structure in the data and allows it to track subtle trends. beck2016 used 100 neighbors in their fitting procedure, though they implemented more nuanced criteria to handle under-sampled regions of color–magnitude space and to iteratively filter redshift outliers.
beck2016 trained their model using 1,976,978 galaxies with spec-’s from the entire SDSS DR12 spectroscopic sample—specifically, the MGS, the luminous red galaxy sample (Eisenstein_LGS), and the Baryon Oscillation Spectroscopic Survey (BOSS; Dawson_BOSS)—plus additional spectroscopy from other surveys, such as DEEP2 (Newman_2013), VIPERS (Garilli_VIPERS; Guzzo_VIPERS), and zCOSMOS (Lilly_2009). beck2016 estimated photo-’s for the 208,474,076 galaxies of the SDSS primary photometric catalog, thus their aims were broader than those of the other approaches we consider, which only used MGS galaxies for training and testing. That being said, MGS galaxies dominate the information content within the color–magnitude–redshift space covered by MGS galaxies, which is further prioritized by the local nature of their regression method.
@@ 3.2 Random Forest
Inspired by previous random forest photo- efforts (e.g., carliles2010RFphotoz; CarrascoKind2013; zhou2019), we trained a random forest (Breiman2001RF) for photo- estimation using the RandomForestRegressor method from the Python package Scikit-learn (PedregosaEtal2011Sklearn) as a classical ML alternative to the local linear regression method of beck2016. Random forest regression is a simple supervised learning algorithm consisting of a “forest” of decision trees that are each trained to predict a value with the final result of the algorithm being the average of each tree’s predicted value. To train the algorithm, we used the training set described in Section 2. We tested providing the random forest with multiple different sets of input features, starting with the features beck2016 used to train their network, as discussed in Section 3.1. To test the impact of input feature choices, we ran random forest models with several combinations of input features:
@@
Ultimately, we found only minor differences in photo- performance (both globally and within individual SOM cells) by adding more features, so we chose model 1 (the same features as beck2016) as our fiducial random forest photo- model. However, we adopt the beck2016 photo-’s as our fiducial classical ML photo- method because it produced qualitatively similar results to the random forest approach, but the beck2016 photo-’s are widely used and are a standard comparison sample for photo- analyses.
@@ 3.3 Deep Learning Models
Despite significant differences in architectures, loss functions, and training mechanisms, a variety of modern deep learning models—including CNNs (pasquet2019), semi-supervised representation learning (hayat2021), and deep capsule networks (dey2022)—have all significantly out-performed classical ML photo- methods (namely beck2016) on the SDSS MGS. In this work, we considered photo-’s from pasquet2019 and dey2022, both of which were trained using the dataset described in Section 2. We adopt the latter as our fiducial deep learning photo- model, though we found very similar results with the former.
@@ 3.3.1 Capsule Networks
Hinton2011CapsNet proposed the idea of capsule networks as an alternative to CNNs, where the artificial neurons of a capsule network are organized into groups (capsules), producing the vector analog to the scalar neurons in CNNs such that they are viewpoint equivariant (as opposed to the viewpoint invariance of CNNs) and therefore could be more efficient to train. However, actually training capsule networks proved challenging until the introduction of the dynamic routing algorithm by SabourEtal2017Capsnet that dramatically improved the viability of their training. Soon after, katebi2019 became the first work to apply capsule networks to galaxy images for galaxy morphology prediction.
dey2022 utilized a deep capsule network for their encapzulate model (compared to the shallow network proposed by SabourEtal2017Capsnet) developed by RajasegaranEtal2019Deepcaps. Its architecture consists of three sub-networks: a deep capsule network for encoding the galaxy images to a low dimensional embedding space, a class-independent decoder network that helps the network generalize with loss terms for image reconstruction and morphological prediction (using disk/spheroid labels from Galaxy Zoo 1; Lintott2011GalaxyZoo1), and a redshift prediction network.
@@ 3.3.2 Global Photo-z Performance
Figure 1 shows versus for the beck2016 and encapzulate photo-’s of SDSS MGS test set galaxies. Compared to the beck2016 photo-’s, the encapzulate photo-’s have a smaller scatter as measured using the normalized median absolute deviation:
where . In the case of photo-’s, we define as (, ). The encapzulate photo-’s also have a significantly lower fraction of catastrophic outliers (as defined by , where ) and a smaller global bias () than the beck2016 photo-’s. These are generic results of comparing classical ML and deep learning photo- methods for the SDSS dataset, though it is not understood how the deep learning models leverage additional redshift information from the images.
Furthermore, there is a slight but perceptible trend of the beck2016 photo-’s being biased towards the mean redshift of the dataset, such that they are overestimated at low and underestimated at high , which is a common bias for ML methods called attenuation bias (see Section 6.1). Below, we compare classical ML and deep learning photo-’s for sub-samples in small neighborhoods in color space to understand how local trends impact global metrics.
@@ 4 Partitioning Color Space with a Self-Organizing Map
To study the impact of attenuation bias in local neighborhoods of color space, we partition color space using a SOM. A SOM is an unsupervised dimensionality reduction algorithm that works by grouping together objects with similar properties into discrete cells, with objects in a cell having properties that are most similar to other objects in the same cell and neighboring cells usually containing galaxies that are broadly similar. Since SOMs produce easily interpretable results with clear visualizations, they have been used for photo- estimation (e.g., CarrascoKindandBrunner2014; Myles2021; Wright2025) and characterizing the coverage of color space by objects with spec-’s (e.g., Masters2015; Masters2017; Stanford2021; Gruen2023; McCullough2024).
For our implementation, we used the minisom python package (vettigli2018minisom) to produce a two-dimensional SOM with 10 12 cells using the , , , and colors of all objects in our sample (i.e., including objects in both the training and test sets). Relative to other SOMs, such as the 75 150 cell SOM used in Masters2015, our SOM more coarsely bins color space. This choice was motivated by the comparably narrow range in color and redshift of our sample and the need to be able to feasibly perform visual inspection of photo- diagnostic plots for each SOM cell.
Figure 2 shows the SOM color-coded by each cell’s mean spec- (left panel) and mean bulge-to-total ratio (B/T; right panel). Generally speaking, the mean spec- increases from left to right and B/T increases from bottom to top.
As representatives of the two main galaxy populations, we chose Cell A (8, 9), which contains red elliptical galaxies (high B/T), and Cell B (1, 3), which contains blue disks (low B/T). The other two cells of interest show more unusual scenarios. Cell C (3, 4) has an intermediate B/T and is located in the transition region between low and high B/T. It contains a diverse collection of objects, including red disk galaxies (Masters2010), post-starburst galaxies (Dressler1983), and galaxies with active galactic nuclei, determined by examining the SDSS and DESI-LS images and SDSS spectra of galaxies in this cell. Finally, Cell D (0, 0) contains very low- galaxies (), a regime where the beck2016 algorithm suffered from a large fraction of catastrophic outliers (geha2017) but deep learning methods performed well (e.g., Wu2022; DarraghFord2023).
@@ 5.1 Photo-z Scatter in Individual SOM Cells
Figure 3 shows the ratio between the values of the beck2016 and encapzulate photo-’s for individual SOM cells (gray points) as a function of the cell’s B/T. For all of the SOM cells, the for the beck2016 photo-’s is at least as large as that of encapzulate with most of the cells having an value – times larger. The highest B/T cells typically have smaller from encapzulate, though a few individual cells (like Cell A) approach parity. The solid blue line and blue shaded regions show the binned median trend and its estimated uncertainty (16th–84th percentile range), respectively, the latter being calculated using a smoothed bootstrap. The first bin (0 B/T 0.1) only contained one SOM cell, so we assumed that the uncertainty on its median value would be the average uncertainty of the other bins multiplied by (i.e., the standard error on the median), where . The binned medians show a statistically significant trend that is slightly decreasing from B/T with an ratio of about –, which is similar to the global ratio of 1.5.
We found similar results when using the random forest photo-’s in lieu of the beck2016 photo-’s and when using the pasquet2019 photo-’s instead of the encapzulate photo-’s. This suggests that the deep learning methods are harnessing additional redshift information from the spatially-resolved images that purely photometry-based ML methods cannot access, particularly for SOM cells with low and intermediate B/T.